# A note on the triple product property for finite groups with abelian normal subgroups of prime index

### v5

@@ Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, and $t,t^{\prime}\in T$, and $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. If this is the case then $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [5, Conjecture 7.6]. This note proves a more general version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group which contains an abelian normal subgroup of prime index $p$, and discusses its implications for $\rho$ for groups with cyclic normal subgroups of prime index, based on the known data for $\rho$ in such groups of small order.
@@ 1. Introduction
(1.1) $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1\Longrightarrow s=s^{\prime},t=t^{\prime},u=u^{\prime},$
@@
TPP triples were first introduced by H. Cohn and C. Umans in 2003 to study the complexity of fast matrix multiplication [3]. Further applications and developments of this approach were described in [2, 5, 6]. However this note pursues the combinatorial problems for TPP triples in finite groups introduced by Peter M. Neumann in [7]. These involve two quantities. The first quantity, called the subgroup TPP ratio of $G$, is defined as
(1.3) $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|},$
@@
As subgroup TPP triples are special cases of TPP triples which are composed only of subgroups, clearly $\beta_{0}(G)\leq\beta(G)$ and $\rho_{0}(G)\leq\rho(G)$. Note that $G$ always realises the trivial subgroup TPP triple $(G,\{1\},\{1\})$, so that $\beta_{0}(G)\geq|G|$, or equivalently, $\rho_{0}(G)\geq 1$.
@@
using the following observation [7, Observation 3.1]:
@@
The author is aware of a proof (from a private communication from another researcher) of this conjecture for dihedral groups $D_{2n}$, i.e. that $\rho(D_{2n})\leq\frac{4}{3}$.
This note focuses on subgroup TPP ratio $\rho_{0}$ (equivalently, subgroup TPP capacity $\beta_{0}$), and proves a generalisation of the conjecture for $\rho_{0}$ for all groups with abelian normal subgroups of prime index. First, a few relevant properties and characterisations of TPP triples are stated, followed by proofs of two technical results needed to establish the main result. Following the proof of the main result some implications for $\rho$ for groups with cyclic normal subgroups of prime index are discussed, based on known data for $\rho$ in such groups.
@@ 2. Elementary Properties of TPP Triples
For non-empty subsets $X,Y\subseteq G$ the quotient set $Q(X,Y)$ of $X$ and $Y$ is the subset $Q(X,Y)=XY^{-1}=\{xy^{-1}|x\in X,y\in Y\}\subseteq G$. If $X=Y$ then $Q(X,Y)=Q(X)=XX^{-1}$.
@@
(2.2) $Q(S)\cap Q(T)Q(U)=Q(T)\cap Q(U)=\{1\},$
And if $S,T,U$ happen to be subgroups then the defining relation above simplifies to
@@
    If $\pi\in S_{3}$, i.e. $\pi$ is a permutation of $\{S,T,U\}$, then $(S^{\pi},T^{\pi},U^{\pi})$ is a TPP triple of $G$ (permutation invariance) [3, Lemma 2.1].
@@
Given a TPP triple $(S,T,U)$ of a group $G$ and a subgroup $H\leq G$ there is a TPP triple $(S\cap H,T\cap H,U\cap H)$ of $H$.
@@
Part (2) of Proposition 2.5 means that only non-abelian groups can realise TPP triples of non-trivial size.
3. Technical Results
Some technical results needed for the main result are stated and proved. These are based on the idea of decomposing a TPP triple of a given group into smaller TPP triples obtained by independently decomposing the members of the triple with the left (or right) cosets of a suitable subgroup, using the set-closure property of TPP triples in relation to subgroups (see Observation 2.4).
@@ 3. Technical Results
Let $G$ be a group, $H\leq G$ and $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$ whose order is denoted by $|G:H|$ (also, $[G:H]$). For a non-empty subset $S\subseteq G$ let the $H$-support of $S$ be the set $\overline{S}:=\{gH\in G/H\mid S\cap gH\neq\emptyset\}\subseteq G/H$, i.e. the set of all cosets $gH\in G/H$ which intersect with $S$. If $(S,T,U)$ is a TPP triple of $G$ let its $H$-restriction be the triple $(S_{0},T_{0},U_{0})$, where $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$, and this is a TPP triple of $H$.
@@
Let $G$ be a group and $H\leq G$.
(1) If $S\leq G$ and $\overline{S}$ is its $H$-support, then $|\overline{S}|=|S:S\cap H|$.
(2) If $S\leq G$ and $H\trianglelefteq G$, i.e. $H$ is normal in $G$, then $\overline{S}\leq G/H$.
(3) If $H$ is abelian and $(S,T,U)$ is a subgroup TPP triple of $G$ then $|S||T||U|\leq\frac{\sigma\tau\upsilon}{n}|G|$, where $|G:H|=n$, and $\sigma=|S:S\cap H|$, $\tau=|T:T\cap H|$, $\upsilon=|U:U\cap H|$.
@@
Let $G$ and $H$ be given as above and $|G:H|=n$.
(1) Let $S\leq G$ and $\overline{S}$ be its $H$-support as defined above. Let $gH\in\overline{S}$ and $S_{g}:=S\cap gH$. By definition $S_{g}\neq\emptyset$. If $x\in S_{g}$ then $|S_{g}|=|S\cap gH|=|xS\cap xH|=|x(S\cap H)|=|S\cap H|$, where $x(S\cap H)$ is a coset of $S\cap H\leq S$. As the cosets $gH\in G/H$ are disjoint so are the sets $S_{g}$, which thus partition $S$. If $\sigma=|\overline{S}|$ then $|S|=\sigma|S\cap H|$, i.e. $\sigma=|S:S\cap H|$. The same result holds if left cosets are replaced by right cosets.
(2) Let $S\leq G$. If $H\trianglelefteq G$ then using the Second Isomorphism Theorem $SH\leq G$ and $\overline{S}=SH/H\leq G/H$, where $SH/H\cong S/(S\cap H)$.
(3) Let $(S,T,U)$ be a subgroup TPP triple of $G$. Then by Observation 2.4 $(S\cap H,T\cap H,U\cap H)$ is a subgroup TPP triple of $H$. If $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$ then by part (1) above $\sigma=|S:S_{0}|$, $\tau=|T:T_{0}|$, $\upsilon=|U:U_{0}|$, and, as by assumption $H$ is abelian, by Proposition 2.5 it follows that $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$, which shows that $|S||T||U|\leq\sigma\tau\upsilon|H|=\frac{\sigma\tau\upsilon}{n}|G|$. ∎
@@
Let $G$ be a group, $H\trianglelefteq G$ an abelian subgroup, $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Furthermore, let the $H$-supports $\overline{S}$, $\overline{T}$, $\overline{U}$ be defined as in Observation 3.2, where these are subgroups of $G/H$.
(1) If $xH,yH$ are any non-trivial cosets of $H$ such that $S_{x},S_{y}\in\overline{S}$, $T_{x}\in\overline{T}$ and $U_{y}\in\overline{U}$, and letting $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, and $U_{0}:=U\cap H$, then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are non-trivial cosets of $S_{0}T_{0}U_{0}$ in $H$, that is, $S_{x}^{-1}T_{x}U_{0},S_{y}^{-1}U_{y}T_{0}\neq S_{0}T_{0}U_{0}$, and also distinct, that is, $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$.
(2) If $xH,yH$ are any non-trivial and distinct cosets of $H$ then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$.
@@
Let $H$ be as given ($H$ is abelian and normal in $G$), and $H<G$ can be assumed, otherwise there is nothing to prove. Several facts used in the proof are noted for convenience:
@@
  • If $x_{1},x_{2}\in xH$ are any coset elements then $x_{1}x_{2}^{-1}\in H$ (if $x_{1}=xh_{1},x_{2}=xh_{2}$ for some $h_{1},h_{2}\in H$ then $x_{1}x_{2}^{-1}=xh_{1}h_{2}^{-1}x^{-1}=xx^{-1}h_{3}=h_{3}\in H$ for some $h_{3}\in H$ (by the normality of $H$ in $G$).
  • Every permutation $(S^{\pi},T^{\pi},U^{\pi})$ of the TPP triple $(S,T,U)$, for $\pi\in S_{3}$, is also a TPP triple (permutation invariance, Observation 2.3).
@@
  • The subgroups $S_{0},T_{0},U_{0}$ as defined above, which satisfy the TPP, also have the property that $S_{0}\trianglelefteq S$, $T_{0}\trianglelefteq T$, $U_{0}\trianglelefteq U$ and $SH/H\cong S/S_{0}$, $TH/H\cong T/T_{0}$, $UH/H\cong U/U_{0}$ (Second Isomorphism Theorem). This means that translates of the form $S_{0}s_{x}=s_{x}S_{0},T_{0}t_{y}=t_{y}T_{0},U_{0}u_{z}=u_{z}U_{0}$, for $s_{x}\in S\cap xH$, $t_{y}\in T\cap yH$, $u_{z}\in U\cap zH$ and $xH,yH,zH\in G/H$, are cosets of $S_{0},T_{0},U_{0}$ respectively in $S,T,U$ respectively which intersect with the cosets $xH,yH,zH$ respectively.
@@
  • As $H$ is abelian, the subgroups $S_{0}$, $T_{0}$, and $U_{0}$, are also normal in $H$ and thus form a set product $S_{0}T_{0}U_{0}$ which is a subgroup of $H$ of order $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$ (using the injective triple product map on $S_{0}\times T_{0}\times U_{0}$ into $H$ from Proposition 2.5).
@@
(1) Let $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$, and $S_{x},S_{y},T_{x},U_{y}$ be defined as above. Note that the sets $S_{x}^{-1}=(S\cap xH)^{-1}=S^{-1}\cap(xH)^{-1}=S\cap x^{-1}H$ and $S_{y}^{-1}=(S\cap yH)^{-1}=S\cap y^{-1}H$ and $S_{x}^{-1}T_{x},S_{y}^{-1}T_{y},S_{y}^{-1}U_{y}\subseteq H$ will also appear by calculation. As $S_{x},T_{x}\neq\emptyset$ elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen. Then, as $s_{x}^{-1}t_{x}\in H$ it follows that $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}\subseteq H$, which shows that $S_{x}^{-1}T_{x}U_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$ of size $|S_{x}^{-1}T_{x}U_{0}|=|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Additionally, if $xH\neq H$ (if $x\in G\backslash H$) then this is coset is non-trivial, i.e. $S_{x}^{-1}T_{x}U_{0}\neq S_{0}T_{0}U_{0}$. To see this, suppose there are elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$, $u_{0}\in U_{0}$ and $s_{0}\in S_{0}$, $t_{0}\in T_{0}$, $\widetilde{u}_{0}\in U_{0}$, such that $s_{x}^{-1}t_{x}u_{0}=s_{0}^{-1}t_{0}\widetilde{u}_{0}$. This can be rearranged as $s_{x}s_{0}^{-1}t_{0}\widetilde{u}_{0}u_{0}^{-1}t_{x}^{-1}=1$. As $t_{0}$ and $\widetilde{u}_{0}u_{0}^{-1}$ commute in $H$ this can be further rearranged as $s_{x}s_{0}^{-1}\widetilde{u}_{0}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $(S,U,T)$ implies that $s_{0}=s_{x}$ and $t_{0}=t_{x}$, a contradiction. By transposing $T$ and $U$, and $x$ and $y$, in the argument above, it can also be shown that $S_{y}^{-1}U_{y}T_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and, furthermore, a non-trivial coset if $yH\neq H$.
Suppose $xH,yH\neq H$ and $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}\neq\emptyset$. Then there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $u_{y}\in U_{y}$, $t_{0}\in T_{0}$, and $u_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}u_{y}t_{0}$, or $s_{x}s_{y}^{-1}u_{y}t_{0}u_{0}^{-1}t_{x}^{-1}=1$. Since $t_{0}$ and $u_{0}$ commute in $H$ this can be rewritten as $s_{x}s_{y}^{-1}u_{y}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $S,U,T$ implies that $u_{y}=u_{0}$ and $t_{x}=t_{0}$, a contradiction. So $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$ if $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$ such that $xH,yH\neq H$, which proves part (1).

### v7

@@ Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, and $t,t^{\prime}\in T$, and $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. If this is the case then $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [5, Conjecture 7.6]. This note proves a more general version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group that contains an abelian normal subgroup of prime index $p$, an improvement by a factor of $\frac{1}{2p-1}$ on the general upper bound of $p^{2}$ when $G$ contains any abelian subgroup of index $p$. In conclusion a generalised conjecture using the same upper bound is presented for $\rho$ for groups with cyclic normal subgroups of prime index, based on the known data for $\rho$ in such groups of small order.
@@ 1. Introduction
(1.1) $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1\Longrightarrow s=s^{\prime},t=t^{\prime},u=u^{\prime}$
@@
TPP triples were first introduced by H. Cohn and C. Umans in 2003 to study the complexity of fast matrix multiplication in the context of finite groups, to be more specific, to realise matrix multiplication as multiplication in a finite group algebra via a triple of non-empty subsets of the group satisfying the TPP, and that are used to index the rows and columns of the matrices being multiplied and then recover the entries of the product [3]. The effectiveness of this group-theoretic approach depends on finding the smallest possible groups that realise TPP triples with parameter types matching the dimensions of matrix multiplication of interest [2, 5, 6].
However this note pursues some combinatorial aspects of maximising TPP triple sizes in finite groups, independently of their applications to matrix multiplication, as described more formally by P. Neumann in [7]. These involve two quantities. The first quantity, called the subgroup TPP ratio of $G$, is defined as
(1.3) $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$
@@
As subgroup TPP triples are special cases of TPP triples that are composed only of subgroups, clearly $\beta_{0}(G)\leq\beta(G)$ and $\rho_{0}(G)\leq\rho(G)$. Note that $G$ always realises the trivial subgroup TPP triple $(G,\{1\},\{1\})$, so that $\beta_{0}(G)\geq|G|$, or equivalently, $\rho_{0}(G)\geq 1$.
@@
using the fact that [7, Observation 3.1]
@@
The author is aware of a proof (from a private communication from another researcher) of this conjecture for dihedral groups $D_{2n}$, that is, $\rho(D_{2n})\leq\frac{4}{3}$.
This note focuses on subgroup TPP ratio $\rho_{0}$ (equivalently, subgroup TPP capacity $\beta_{0}$), and proves a generalisation of the conjecture for $\rho_{0}$ for all groups with abelian normal subgroups of prime index. First, a few relevant properties and characterisations of TPP triples are stated, followed by proofs of two technical results needed to establish the main result. Following the proof of the main result and a basic corollary some implications for $\rho$ for groups with cyclic normal subgroups of prime index are discussed, based on known data for $\rho$ in such groups.
@@ 2. Elementary Properties of TPP Triples
The TPP can also be described in terms of certain sets called quotient sets, which have the following definition.
(2.2) $Q(X,Y)=XY^{-1}=\{xy^{-1}|x\in X,y\in Y\}\subseteq G,\hskip 30.00005ptX,Y\subseteq G;X,Y\neq\emptyset$
If $X=Y$ then $Q(X)$ is a shorthand for $Q(X,Y)$.
Note that, by definition, a quotient set $Q(X)$ contains the identity $1$, is equal to its inverse, that is, $Q(X)=Q(X)^{-1}$, and, furthermore, if $1\in X$ then $Q(X)$ contains both $X$ and its inverse $X^{-1}$, that is, $X\cup X^{-1}\subseteq Q(X)$. However, a quotient set is not necessarily closed under taking products, as otherwise it would be a (sub)group, which isn’t generally true. For a subgroup it is true that it is equal to its quotient set.
@@
(2.3) $Q(S)\cap Q(T)Q(U)=Q(T)\cap Q(U)=\{1\}.$
If $S,T,U$ happen to be subgroups then the defining relation above simplifies to
@@
    If $\pi\in S_{3}$, that is, $\pi$ is a permutation of $\{S,T,U\}$, then $(S^{\pi},T^{\pi},U^{\pi})$ is a TPP triple of $G$ (permutation invariance) [3, Lemma 2.1].
@@
Given a TPP triple $(S,T,U)$ of a group $G$ and a subgroup $H\leq G$ every triple of the form $(S^{\prime}\cap H,T^{\prime}\cap H,U^{\prime}\cap H)$ for non-empty subsets $S^{\prime}\subseteq S$, $T^{\prime}\subseteq T$, $U^{\prime}\subseteq U$, with $1\in S^{\prime}\cap T^{\prime}\cap U^{\prime}$, is a TPP triple of $H$.
Proof.
If $G$ and $H$ are as given, and $(S,T,U)$ is a TPP triple of $G$, then for any non-empty subsets $S^{\prime}\subseteq S$, $T^{\prime}\subseteq T$, $U^{\prime}\subseteq U$, with $1\in S^{\prime}\cap T^{\prime}\cap U^{\prime}$, define the subsets $S^{\prime}_{0}:=S^{\prime}\cap H$, $T^{\prime}_{0}:=T^{\prime}\cap H$, $U^{\prime}_{0}:=U^{\prime}\cap H$. Then $1\in S^{\prime}_{0}\cap T^{\prime}_{0}\cap U^{\prime}_{0}$ and $Q(S^{\prime}_{0})\subseteq Q(S)$, $Q(T^{\prime}_{0})\subseteq Q(T)$, $Q(U^{\prime}_{0})\subseteq Q(U)$, and $Q(S^{\prime}_{0})\cap Q(T^{\prime}_{0})Q(U^{\prime}_{0})=Q(T^{\prime}_{0})\cap Q(U^{\prime}_{0})=\{1\}$. ∎
@@
Part (2) of Proposition 2.5 means that only nonabelian groups can realise TPP triples of non-trivial size.
3. Coset Decomposition of TPP Triples
Some technical results needed for the main result are stated and proved. These are based on the idea of decomposing a TPP triple of a given group into smaller TPP triples obtained by independently decomposing the members of the triple with the left (or right) cosets of a suitable subgroup of small index, as first described by P. Neumann in [7, Observation 4.1].
@@ 3. Coset Decomposition of TPP Triples
Let $G$ be a group, $H$ a subgroup, $G/H=\{gH\mid g\in G\}$ the collection of all (left) cosets of $H$ in $G$, and $|G:H|$ the index (or size) of $G/H$. For a non-empty subset $S\subseteq G$ let the $H$-support of $S$ be the set $\overline{S}:=\{gH\in G/H\mid S\cap gH\neq\emptyset\}\subseteq G/H$, that is, the set of all cosets of $H$ that intersect with $S$. By definition $gH\in\overline{S}\iff S_{g}:=S\cap gH\neq\emptyset$ for any coset $gH\in G/H$. If $(S,T,U)$ is a TPP triple of $G$ let its restriction to $H$, or, simply, $H$-restriction, be the triple $(S_{0},T_{0},U_{0})$, where $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$, and this is a TPP triple of $H$.
@@
Let $G$ be a group and $H$ a subgroup.
(1) If $S$ is a subgroup of $G$ and $\overline{S}$ is its $H$-support then $|\overline{S}|=|S:S\cap H|$.
(2) If $S$ is a subgroup of $G$ and $H$ is normal in $G$ then $\overline{S}$ is a subgroup of $G/H$.
(3) If $H$ is abelian and normal in $G$, and $(S,T,U)$ is a subgroup TPP triple of $G$ then
(3.1) $|S||T||U|\leq\frac{\sigma\tau\upsilon}{n}|G|$
where $n=|G:H|$, and $\sigma=|S:S\cap H|$, $\tau=|T:T\cap H|$, $\upsilon=|U:U\cap H|$.
@@
Let $G$ and $H$ be given as above and $n=|G:H|$.
(1) Let $S\leq G$ and $\overline{S}$ be its $H$-support as defined above. Let $gH\in\overline{S}$ and $S_{g}:=S\cap gH$. By definition $S_{g}\neq\emptyset$. If $x\in S_{g}$ then $|S_{g}|=|S\cap gH|=|xS\cap xH|=|x(S\cap H)|=|S\cap H|$, where $x(S\cap H)$ is a coset of $S\cap H\leq S$. As the cosets $gH\in G/H$ are disjoint so are the sets $S_{g}$, which thus form an equal-sized partition of $S$. If $\sigma=|\overline{S}|$ then $|S|=\sigma|S\cap H|$, i.e. $\sigma=|S:S\cap H|$. The same result holds if left cosets are replaced by right cosets.
(2) Let $S\leq G$. If $H\trianglelefteq G$ then $G/H$ is a (quotient) group and using the Second Isomorphism Theorem $SH\leq G$ and $\overline{S}=SH/H\leq G/H$, where $SH/H\cong S/(S\cap H)$.
(3) Let $H\trianglelefteq G$ be abelian, and $(S,T,U)$ be a subgroup TPP triple of $G$. Then by Observation 2.4 $(S\cap H,T\cap H,U\cap H)$ is a subgroup TPP triple of $H$. If $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$, and $\overline{S},\overline{T},\overline{U}$ are the $H$-supports of $S,T,U$ respectively, then by part (1) above $\sigma=|\overline{S}|=|S:S_{0}|$, $\tau=|\overline{T}|=|T:T_{0}|$, $\upsilon=|\overline{U}|=|U:U_{0}|$, where $\overline{S},\overline{T},\overline{U}\leq G/H$. Also, by assumption $H$ is abelian, so by Proposition 2.5 it follows that $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$, which shows that $|S||T||U|\leq\sigma\tau\upsilon|H|=\frac{\sigma\tau\upsilon}{n}|G|$. Note here that $\sigma,\tau,\upsilon$ are divisors of $n$, and of $|S|,|T|,|U|$ respectively, so $\sigma\leq\text{min}\{n,|S|\}$, $\tau\leq\text{min}\{n,|T|\}$, $\upsilon\leq\text{min}\{n,|U|\}$. ∎
Note that (3.1) can also be derived for the more general case of subset TPP triples, leading to the general upper bound
(3.2) $|S||T||U|\leq n^{2}|G|$
for the size of any TPP triple $(S,T,U)$ of a group $G$ with an abelian subgroup of index $n$ [7, Corollary 4.2].
@@
Let $G$ be a group, $H$ an abelian normal subgroup, $G/H$ the quotient group, and $(S,T,U)$ a subgroup TPP triple of $G$. Furthermore, let the $H$-supports $\overline{S}$, $\overline{T}$, $\overline{U}$ be defined as in Observation 3.2, where these are subgroups of $G/H$, the subgroups $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$, and the sets $S_{x}:=S\cap xH$, $T_{y}:=T\cap yH$, $U_{z}:=U\cap zH$ for cosets $xH,yH,zH\in G/H$.
(1) If $xH\in\overline{S}\cap\overline{T}$, then the set $S_{x}^{-1}T_{x}U_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and, moreover, non-trivial if $xH$ is non-trivial.
(2) If $xH,yH\in\overline{S}\cap\overline{T}$ are distinct, and $U_{0}$ is normal in $G$, then the cosets $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}T_{y}U_{0}$ are distinct.
(3) If $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$, and at least one of them is non-trivial, then the cosets $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are distinct.
@@
Let $H$ be as given ($H$ is abelian and normal in $G$). Several basic facts used in the proof are recalled for convenience:
@@
  • If $x_{1},x_{2}\in xH$ are any coset elements then $x_{1}x_{2}^{-1}\in H$ (if $x_{1}=xh_{1}$ and $x_{2}=xh_{2}$ for some $h_{1},h_{2}\in H$ then $x_{1}x_{2}^{-1}=xh_{1}h_{2}^{-1}x^{-1}=xx^{-1}h_{3}=h_{3}\in H$ for some $h_{3}\in H$ (by the normality of $H$ in $G$).
  • Every permutation $(S^{\pi},T^{\pi},U^{\pi})$ of the TPP triple $(S,T,U)$ of $G$, for $\pi\in S_{3}$, is also a TPP triple of $G$ (permutation invariance, Observation 2.3).
@@
  • The subgroups $S_{0},T_{0},U_{0}$ as defined above, which satisfy the TPP, are normal in $S,T,U$ respectively, and $SH/H\cong S/S_{0}$, $TH/H\cong T/T_{0}$, $UH/H\cong U/U_{0}$ (Second Isomorphism Theorem). This means that translates of the form $S_{0}s_{x}=s_{x}S_{0},T_{0}t_{y}=t_{y}T_{0},U_{0}u_{z}=u_{z}U_{0}$, for elements $s_{x}\in S\cap xH$, $t_{y}\in T\cap yH$, $u_{z}\in U\cap zH$ and cosets $xH,yH,zH\in G/H$, are cosets of $S_{0},T_{0},U_{0}$ respectively in $S,T,U$ respectively that intersect with the cosets $xH,yH,zH$ respectively, and contain the elements $s_{x},t_{y},u_{z}$ respectively, that is, $s_{x}S_{0}=S_{0}s_{x}=S_{x}$, $t_{y}T_{0}=T_{0}t_{y}=T_{y}$, $u_{z}U_{0}=U_{0}u_{z}=U_{z}$.
@@
  • As $H$ is abelian, the subgroups $S_{0}$, $T_{0}$, and $U_{0}$, are also normal in $H$ and thus form a set product $S_{0}T_{0}U_{0}$ that is a subgroup of $H$ of order $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$ (using the injective triple product map on $S_{0}\times T_{0}\times U_{0}$ into $H$ from Proposition 2.5).
@@
(1) Let $xH\in\overline{S}\cap\overline{T}$. Then, by definition, the sets $S_{x},T_{x}$ are non-empty, and elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen. As $s_{x}^{-1}t_{x}\in H$ a translate of $S_{0}T_{0}U_{0}$ in $H$ can be formed, which is $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}$ and thus a coset of $S_{0}T_{0}U_{0}$ in $H$. Its size is $|S_{x}^{-1}T_{x}U_{0}|=|s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})|=|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. If $xH\neq H$ (if $x\in G\backslash H$) then the coset $S_{x}^{-1}T_{x}U_{0}$ is non-trivial, that is, $S_{x}^{-1}T_{x}U_{0}\cap S_{0}T_{0}U_{0}=\emptyset$ if $xH\neq H$. To see this, suppose there are elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$, $u_{0}\in U_{0}$ and $s_{0}\in S_{0}$, $t_{0}\in T_{0}$, $\widetilde{u}_{0}\in U_{0}$, such that $s_{x}^{-1}t_{x}u_{0}=s_{0}^{-1}t_{0}\widetilde{u}_{0}$. This can be rearranged as $s_{x}s_{0}^{-1}t_{0}\widetilde{u}_{0}u_{0}^{-1}t_{x}^{-1}=1$. As $t_{0}$ and $\widetilde{u}_{0}u_{0}^{-1}$ commute in $H$ this can be further rearranged as $s_{x}s_{0}^{-1}\widetilde{u}_{0}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $(S,U,T)$ implies that $s_{0}=s_{x}$ and $t_{0}=t_{x}$, which is a contradiction if $xH\neq H$.
Note here that by transposing $T$ and $U$ it can be shown that there are cosets of the form $s_{x}^{-1}u_{x}(S_{0}T_{0}U_{0})=s_{x}^{-1}u_{x}(S_{0}U_{0}T_{0})=S_{x}^{-1}U_{x}T_{0}$ in $H$, and which are non-trivial if $xH$ is non-trivial.